The following regression model was fitted to sample data with 12 observations: y Overscript ̂ EndScripts= 30 + 4.50x. What is the residual for an observation (x = 2, y = 40)?
Answer #1

Solution :

Predicted = 30 + 4.50 * 2 = 39

Residual = Actual - Predicted

= 40 - 39

= 1